# Antifragile Perimeter Control: Anticipating and Gaining from Disruptions with Reinforcement Learning

### v4

@@ 1. General notations in problem formulation
$M_{ij}(t)$ Trip completion with OD from region $i$ to $j$ $(i\neq j)$ at time $t$
@@
2. Notations in reinforcement learning
@@
$s_{t}$ $s_{t}\in\mathcal{S}$, the observable state in simulation at time $t$
@@
$a_{t}$ $a_{t}\in\mathcal{A}$, the action taken in simulation at time $t$
@@
$r_{t}$ $r_{t}=\mathcal{R}(s_{t},a_{t})$, the received reward with state $s_{t}$ and action $a_{t}$ at time $t$
@@
$Q(s_{t},a_{t})$ Expected long-term return for taking action $a_{t}$ in state $s_{t}$ at time $t$
@@
$\rho^{\beta}$ All possible trajectories of $s_{t}$
$I$ The objective function for the actor-network
3. Notations for the additional antifragile terms applied in reinforcement learning
$r_{com}(t)$ Reward term in RL based on trip completion, equaling to $\sum^{2}_{i=1}M_{ii}(t)$
$r_{red}(t)$ Additional reward term in RL based on derivatives and redundancy
$r_{dam}(t)$ Additional damping term to counter possible action oscillation
$\omega_{h}$ The weight of first derivative in the additional reward term $r_{red}(t)$
$\omega_{\Delta h}$ The weight of second derivative in the additional reward term $r_{red}(t)$
$\alpha_{i}(t)$ Binary variable determining the term to be reward/penalty
$h_{i}(t)$ The first derivative of traffic state at time $t$
$\Delta h_{i}(t)$ The second derivative of traffic state at time $t$
Eq. 1a describes the change rate of the intraregional vehicle accumulation of region $i$. It is the sum of intraregional traffic demand in this region $q_{ii}(t)$ with the perimeter control regulated transfer flow $u_{ji}(t)\cdot M_{ji}(t)$, then deducted by trip completion $M_{ii}(t)$. similarly, the change rate of interregional traffic accumulation, as Eq. 1b shows, is the difference between the interregional traffic demand $q_{ij}(t)$ and the regulated transfer flow $u_{ij}(t)\cdot M_{ij}(t)$:
@@ 3. Notations for the additional antifragile terms applied in reinforcement learning
$\frac{dn_{ij}(t)}{dt}$ $=q_{ij}(t)-u_{ij}(t)\cdot M_{ij}(t),\>(i\neq j)$ (1b)
The total trip completion $M_{i}(t)$ for region $i$ at time $t$ is calculated based on the trip accumulation and the related MFD $G_{i}(n_{i}(t))$, which is the sum of the intraregional trip completion $M_{ii}(t)$ and the interregional transfer flow $M_{ij}(t)\>(i\neq j)$:
@@
Intraregional and interregional vehicle accumulation $n_{ii}(t)$ and $n_{ij}(t)$ are non-negative values, and $n_{i,\rm cap}$ is the maximal possible number of vehicles accumulated in region $i$, at which value a gridlock will occur in the network. $u_{\min}$ and $u_{\max}$ represent the lower and upper limits for the perimeter control variable $u_{ij}(t)$ for both directions. Such boundaries are in line with Geroliminis et al. (2013); Zhou and Gayah (2021), as perimeter control is normally implemented through signalization. While $u_{\max}$ accounts for the lost time between the red and green phases, $u_{\min}$ is essential since an indefinitely long red light is rare in real-world urban networks.
In contrast to the objective function to be applied for control-based strategies, the reward function for the proposed antifragile RL-based algorithms $J_{anti}$ is in Eq. 4, with $r_{com}(t)$ standing for the same trip completion as in Eq. 3, i.e., $r_{com}(t)=\sum_{i=1,2}M_{ii}(t)$.
[…]
The second term $r_{dam}(t)$ is the damping term to counter the oscillating behavior of the RL agents due to the incorporation of derivatives in RL. The third term $r_{red}(t)$ represents the redundancy term that aims to build up a proper redundancy so that the proposed RL algorithm does not reward the agent for targeting the optimal critical accumulation. More explanation of the damping term $r_{dam}(t)$ and the redundancy term $r_{red}(t)$ can be found in Section 4.3.
@@ 4.1 Model predictive control and moving horizon estimation
Model Predictive Control (MPC) is a widely recognized control approach extensively used for regulating dynamic systems across various engineering fields, particularly in transportation and perimeter control. Readers interested in MPC and its applications in traffic engineering can refer to Geroliminis et al. (2013); Haddad and Mirkin (2017). The MPC toolkit applied in this paper is introduced in Lucia et al. (2017), which uses the CasADi framework (Andersson et al., 2019) and the NLP solver IPOPT (Wächter and Biegler, 2006).
@@
In Sirmatel and Geroliminis (2020), Moving Horizon Estimation (MHE) is introduced to form an MPC-MHE framework for perimeter control. MHE enables real-time estimation of internal system states by solving an optimization problem that minimizes the discrepancy between predicted outputs and noisy sensor measurements. It further addresses the data observability issues in Geroliminis et al. (2013) as illustrated in Fig. 1, that is, the information on both the OD accumulation $n_{ij}(t)$ and the newly generated demand $q_{ij}(t)$ is almost impossible to acquire in real-time. Therefore, four different combinations of data availability have been tested, i.e., one from either $n_{ij}(t)$ or $[n_{i}(t),M_{ij}(t)]$ and another one from either $q_{ij}(t)$ or $q_{i}(t)$. The accumulation $n_{i}(t)$ for region $i$ can be approximated with loop detector measurements, while interregional trip completion $M_{ij}(t)$ can be directly counted through detectors at the intersections forming the perimeter border. The authors also claim $q_{i}(t)$ can be available with some applications, such as a substantial percentage of GPS information being readily accessible. However, due to the nature of the prediction horizon in MPC-MHE, such data has to be made available in both real-time and the near future, especially under disruptive conditions, rendering it impossible to collect. We will further tackle such observability issues with the proposed antifragile perimeter control algorithm in the following Section 4.3.
4.2 RL algorithm
@@
In RL algorithms, an agent or multiple agents interact with a preset environment and improve its decision-making capacilities, defined as action $a_{t}$ in an action space $\mathcal{A}$, based on the observable state $s_{t}$ in the state space $\mathcal{S}$ and the reward, defined as $r_{t}=\mathcal{R}(s_{t},a_{t})$, where $\mathcal{R}$ is the reward function. The improvement of decision-making is commonly realized through a deep neural network as a function approximator. The RL algorithm applied in this work is Deep Deterministic Policy Gradient (DDPG), as proposed in Lillicrap et al. (2015). By applying an actor-critic scheme, DDPG can manage a continuous action space instead of a discrete action space. Furthermore, Zhou and Gayah (2021) has demonstrated that an RL algorithm with a continuous action space can achieve superior performance. The DDPG algorithm can be divided into two main components, namely the actor and the critic, which are updated at each step through policy gradient and Q-value, respectively. The scheme of the DDPG algorithm applied in this paper is schematically illustrated in Fig. 3.
Figure 3: DDPG scheme
@@
The actor-network is represented by $\mu(\cdot)$, and it determines the best action $a_{t}$ for the perimeter controller based on the current state $s_{t}$. The critic network $Q(\cdot)$ is responsible for evaluating whether a specific state-action pair at time step $t$ yields the maximal possible discounted future reward $Q(s_{t},a_{t})$. A common technique used in DDPG is to create a target actor network $\mu^{\prime}(\cdot)$ and a target critic network $Q^{\prime}(\cdot)$, which are a copy of the original actor and critic network but updated posteriorly to stabilize the training process and prevent overfitting (Zhang et al., 2021), with the target maximal discounted future reward calculated as:
$y_{i}=r_{i}+\gamma Q^{\prime}(s_{i+1},\mu^{\prime}(s_{i+1}|\theta^{\mu^{\prime}})|\theta^{Q^{\prime}})$ (5)
@@
The critic network can then be updated by calculating the temporal difference between the predicted reward and the target reward and minimizing the loss for a mini-batch $N$ sampled from the replay buffer:
$L=\frac{1}{N}\sum_{i}(y_{i}-Q(s_{i},a_{i}|\theta^{Q}))^{2}$ (6)
@@
$I=$ $\mathbb{E}_{s_{t}\sim\rho^{\beta}}[r(s,\mu(s|\theta^{\mu}))|_{s=s_{t}}]$ (7a)
$\nabla_{\theta^{\mu}}I=$ $\mathbb{E}_{s_{t}\sim\rho^{\beta}}[\nabla_{a}Q(s,a|\theta^{Q})|_{s=s_{t},a=\mu(s_{t})}\nabla_{\theta_{\mu}}\mu(s|\theta^{\mu})|_{s=s_{t}}]$ (7b)
$\nabla_{\theta^{\mu}}I\approx$ $\frac{1}{N}\sum_{t}\nabla_{a}Q(s,a|\theta^{Q})|_{s=s_{t},a=\mu(s_{t})}\nabla_{\theta^{\mu}}\mu(s|\theta^{\mu})|_{s_{t}}$ (7c)
First developed in Horgan et al. (2018), a similar Ape-X architecture as applied in Zhou and Gayah (2021) is also adopted in this work, which allows for multiple generations of simulations in the training process to be included and centrally learned for the best policy. A decaying Gaussian noise is also introduced to encourage the DDPG agent to explore the state space.
When training the agent, we add disruptions into the simulation environment starting from a certain training episode, such as surging traffic demand or MFD disruption. These various forms of volatility ought to optimally elicit various learning and decision-making processes. The RL agent would include the results of their prior choices to create and update their weight parameters in a situation with high outcome volatility and be capable of generating and updating expectations after sensing a change in a high-volatility environment.
@@ 4.3 Antifragility and the antifragile terms in RL
As reviewed in Section. 2.2, proper information can be integrated into the RL algorithms to induce robustness and resilience. Following the same idea, we incorporate antifragile derivatives and redundancy terms based on the methodologies reviewed in Section 2.3. First, the benchmark work in Zhou and Gayah (2021) establishes the state as $s_{t}=[n_{ij}(t),q_{ij}(t)]$, with $q_{ij}(t)$ regarded as an estimate of average daily traffic demand. However, real-time traffic demand $q_{ij}(t)$ can hardly be acquired, particularly in an environment with disruptions. Therefore, we propose a state representation incorporating both the first and second order derivatives of vehicle accumulation, which can be computed as the first and second order differences between the vehicle accumulation $[\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ between two consecutive time steps in the discrete format. Similar to Sirmatel and Geroliminis (2020) discussed in Section 4.1, we also tackle the data observability issue in this work, by integrating trip completion $M_{ij}(t)$ and replacing OD-pair accumulation $n_{ij}(t)$ with simply regional accumulation $n_{i}(t)$ so as its first and second differences. It should be noted that $M_{ij}(t)$ in Sirmatel and Geroliminis (2020) refers specifically to the interregional transfer flow $M_{ij}(t),(i\neq j)$. However, when considering a rather homogenous average trip length, the regional trip completion rate $M_{i}(t)$ is linearly related to the network average flow and can be approximated with ease, as illustrated in (Geroliminis and Daganzo, 2008; Zhou and Gayah, 2021). Then the intraregional completion rate can be further computed as $M_{ii}(t)=M_{i}(t)-M_{ij}(t),(i\neq j)$. For fair comparison, we offer two sets of states for testing under either idealized full observability or real-world limited observability.
@@
The action $a_{t}\in\mathcal{A}$ is defined the same as the control variables $u_{ij}(t)$ in Section 3. For the reward term $r_{t}$, while Zhou and Gayah (2021) uses merely the completion rate, in our proposed algorithm, the reward is defined with additional two terms, i.e., the damping term $r_{dam}(t)$ and the redundancy term $r_{red}(t)$, as Eq. 4 shows. When $[\Delta n_{ij}(t),\Delta^{2}n_{ij}(t)]$ or their $[\Delta n_{i}(t),\Delta^{2}n_{i}(t)]$ are incorporated into the state space $\mathcal{S}$ of the algorithm, significant oscillations in the perimeter control variables can often be observed, especially under scenarios with disruptions. As perimeter control is composed of coordinated traffic lights at the border of different regions, oscillating actions will result in fast varying green splits between consecutive cycles. Even though such oscillations may have a minor impact on traffic performance, the operation of signal lights in the real world should be as stable as possible. Therefore, a damping term $r_{dam}(t)$ is introduced into the reward function to penalize potential oscillatory actions. While the absolute difference between two consecutive actions is always below $1$, the constant $\xi_{1}$ bounds the maximal penalty while the exponent $\xi_{2}$ determines how fast the penalty decays when the change of control variables becomes small, indicating a large $\xi_{2}$ only penalizes the agent when the oscillation is substantial.
$r_{dam}(t)=-\xi_{1}\sum_{i,j\in\{1,2\}}|u_{i,j}(t)-u_{i,j}(t-1)|^{\xi_{2}},\quad i\neq j$ (10)
@@
The last term $r_{red}(t)$ in the objective function $J_{anti}$ in Eq. 4 acts as an additional term to build up redundancy in the system, resembling the buffer times in train scheduling (Corman et al., 2018). However, instead of a fixed buffer time commonly seen among most European railway networks (Jovanović et al., 2017), we create a dynamic redundancy term emphasizing the derivatives in a similar way as the differences of the regional accumulation, considering real-world limited observability. Here, we summarize $r_{red}(t)$ as the sum of two terms $r_{red}(t)=H(t)+\Delta H(t)$, with $H(t)$ being an overall term representing the first derivative and $\Delta H(t)$ representing the second derivative, which can be further expanded as:
$H(t)=$ $\sum_{i=1,2}H_{i}(t)=\sum_{i=1,2}\omega_{h}\cdot h_{i}(t)\cdot\alpha_{i}(t)\cdot f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})$ (11a)
$\Delta H(t)=$ $\sum_{i=1,2}\Delta H_{i}(t)=\sum_{i=1,2}\omega_{\Delta h}\cdot\Delta h_{i}(t)\cdot f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})$ (11b)
@@
Since normalization is common practice in RL to avoid gradient explosion, $\omega_{h}$ and $\omega_{\Delta h}$ are introduced as the weight constants for the derivatives to regulate their impact on the reward $\mathcal{R}$. $h_{i}(t)$ and $\Delta h_{i}(t)$ are the first and second difference of the traffic states on the MFD, $h_{i}(t)$ is defined as the difference of trip completion over vehicle accumulation at the end of a time step versus at the beginning of the same time step, and the second derivative $\Delta h_{i}(t)$ is calculated as the difference between the first differences of two consecutive time steps:
$h_{i}(t)=\frac{M_{i}(t)\ -\ M_{i}(t-1)}{n_{i}(t)\ -\ n_{i}(t-1)}$ (12a)
$\Delta h_{i}\left(t\right)=h_{i}\left(t\right)-h_{i}\left(t-1\right)$ (12b)
@@
The binary variable $\alpha_{i}(t)$ in $H(t)$ was designed to reward the agent when moving towards the desired direction on the MFD. For instance, when the traffic state lies in the congested zone of the MFD, the gradient of any step will be negative. However, a penalty should be applied when vehicle accumulation gets larger, while rewarding the agent when accumulation decreases.
$\alpha_{i}(t)=\begin{cases}1,&\text{if }n_{i}(t)\geq n_{i}(t-1),\\ -1,&\text{otherwise.}\end{cases}$ (13)
@@
Here, $f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})$ computes a reduction factor to constrain the impact of the $r_{red}(t)$ term when the accumulation is around the critical accumulation $n_{i,\rm crit}$, where the $r_{red}(t)$ term should have the greatest impact. A modified trigonometric function is applied to realize this purpose. Other functions, such as normal distribution, should also be valid for achieving the same goal.
$f(n_{i}(t),n_{i,\rm crit},n_{i,\rm cap})=\begin{cases}{\Big{(}1+\cos\Big{(}-\pi\cdot\cfrac{n_{i,\rm crit}-n_{i}(t)}{n_{i,\rm crit}}\Big{)}\Big{)}}/2,&\text{if }n_{i}(t)\geq n_{i,\rm crit},\\ {\Big{(}1+\cos\Big{(}-\pi\cdot\cfrac{n_{i}(t)-n_{i,\rm crit}}{n_{i,\rm cap}-n_{i,\rm crit}}\Big{)}\Big{)}/2},&\text{otherwise.}\end{cases}$ (14)
@@
We illustrate the impact of $H(t)$ and $\Delta H(t)$ on building redundancy on the MFD in Fig. 4 and 5. The first derivative $H(t)$ in Fig. 4(a) rewards the agent when moving toward the critical accumulation $n_{i,\rm crit}$ to maximize trip completion. However, when $n$ approaches $n_{i,\rm crit}$, this term drops significantly and becomes a penalty when $n$ exceeds $n_{i,\rm crit}$. In Fig. 4, we showcase the influence of this term on the MFD. With increasing weight coefficient $\omega_{h}$, $n_{i,\rm crit}$ of the modified MFD becomes lower compared to the original, and the reward decreases faster after the accumulation exceeds $n_{i,\rm crit}$. In this way, redundant overcompensation has been established to prevent accumulation from exceeding the critical accumulation when disruption happens unexpectedly.

### v5

@@ 1. Notations for the problem formulation (continuous form)
$M_{ij}(t)$ Trip completion with OD from region $i$ to $j$ at time $t$
@@
2. Notations in RL (discrete form)
$k$ Current time step, $k=\lfloor t/\Delta t\rfloor$
$K$ Total time steps, $K=T/\Delta t$
@@ 2. Notations in RL (discrete form)
$s_{k}$ $s_{k}\in\mathcal{S}$, the observable state in simulation at timestep $k$
@@
$a_{k}$ $a_{k}\in\mathcal{A}$, the action taken in simulation at timestep $k$
@@
$r_{k}$ $r_{k}=\mathcal{R}(s_{k},a_{k})$, the received reward with state $s_{k}$ and action $a_{k}$ at timestep $k$
@@
$Q(s_{k},a_{k})$ Expected long-term return for taking action $a_{k}$ in state $s_{k}$ at timestep $k$
@@
$\rho^{\beta}$ All possible trajectories of $s_{k}$
II The objective function for the actor-network
3. Notations for the antifragile terms in RL (discrete form)
$r_{com,k}$ Reward term in RL based on trip completion
$r_{red,k}$ Additional reward term in RL based on derivatives and redundancy
$r_{dam,k}$ Additional damping term to counter possible action oscillation
$\omega_{h}$ The weight of the first derivative in the additional reward term $r_{red,k}$
$\omega_{\Delta h}$ The weight of the second derivative in the additional reward term $r_{red,k}$
$\alpha_{i,k}$ Binary variable determining the term to be a reward/penalty
$h_{i,k}$ The first derivative of traffic state at timestep $k$
$\Delta h_{i,k}$ The second derivative of traffic state at timestep $k$
Eq. 1a describes the change rate of the intraregional vehicle accumulation of region $i$. It is the sum of intraregional traffic demand in this region $q_{ii}(t)$ with the perimeter control regulated transfer flow $u_{ji}(t)\cdot M_{ji}(t)$, then deducted by trip completion $M_{ii}(t)$. Likewise, the change rate of interregional traffic accumulation, as Eq. 1b shows, is the difference between the interregional traffic demand $q_{ij}(t)$ and the regulated transfer flow $u_{ij}(t)\cdot M_{ij}(t)$:
@@ 3. Notations for the antifragile terms in RL (discrete form)
$\frac{dn_{ij}(t)}{dt}$ $=q_{ij}(t)-u_{ij}(t)\cdot M_{ij}(t)$ (1b)
The total trip completion $M_{i}(t)$ for region $i$ at time $t$ is calculated based on the trip accumulation and the related MFD $G_{i}(n_{i}(t))$, which is the sum of the intraregional trip completion $M_{ii}(t)$ and the interregional transfer flow $M_{ij}(t)$:
@@
Intraregional and interregional vehicle accumulation $n_{ii}(t)$ and $n_{ij}(t)$ are non-negative values, and $n_{i,\rm cap}$ is the maximal possible number of vehicles accumulated in region $i$, at which value a gridlock will occur in the network. $u_{\min}$ and $u_{\max}$ in Eq. 3c represent the lower and upper limits for the perimeter control variable $u_{ij}(t)$ for both directions. Such boundaries are in line with Geroliminis et al. (2013); Zhou and Gayah (2021), as perimeter control is normally implemented through signalization. While $u_{\max}$ accounts for the lost time between the red and green phases, $u_{\min}$ is essential since an indefinitely long red light is rare in real-world urban networks.
In contrast to the objective function to be applied for control-based strategies, the reward function for the proposed antifragile RL-based algorithms $J_{af}$ is described in Eq. 4 in discrete form, with $r_{com,k}$ standing for the total trip completion in both regions at timestep $k$, that is $r_{com,k}=\sum\limits^{k\Delta t+\Delta t}_{t=k\Delta t}\sum\limits_{i=1,2}M_{ii}(t)$.
$J_{af}=\max\limits_{u_{ij}(k)}{\sum_{k=1}^{K}\left(r_{com,k}+r_{dam,k}+r_{red,k}\right)}$ (4)
The second term $r_{dam,k}$ is the damping term to counter the oscillating behavior of the RL agents due to the incorporation of derivatives in RL. The third term $r_{red,k}$ represents the redundancy term that aims to build up a proper redundancy so that the proposed RL algorithm does not reward the agent for targeting the optimal critical accumulation. More explanation of the damping term $r_{dam,k}$ and the redundancy term $r_{red,k}$ can be found in Section 4.3.
@@ 4.1 Model predictive control and moving horizon estimation
Model Predictive Control (MPC) is a widely recognized control approach extensively used for regulating dynamic systems across various engineering fields, particularly in transportation and perimeter control. Readers interested in MPC and its applications in traffic engineering can refer to Geroliminis et al. (2013). The MPC toolkit applied in this paper is introduced in Lucia et al. (2017), which uses the CasADi framework (Andersson et al., 2019).
@@
In Sirmatel and Geroliminis (2020), Moving Horizon Estimation (MHE) is introduced to form an MPC-MHE framework for perimeter control. MHE enables real-time estimation of internal system states by solving an optimization problem that minimizes the discrepancy between predicted outputs and noisy sensor measurements. It further addresses the data observability issues in Geroliminis et al. (2013) as illustrated in Fig. 1, that is, the information on both the OD accumulation $n_{ij}(t)$ and the newly generated demand $q_{ij}(t)$ is almost impossible to acquire in real-time. Therefore, four different combinations of data availability have been tested, i.e., one from either $n_{ij}(t)$ or $n_{i}(t),M_{ij}(t)]$ and another one from either $q_{ij}(t)$ or $q_{i}(t)$. The accumulation $n_{i}(t)$ for region $i$ can be approximated with loop detector measurements, while interregional trip completion $M_{ij}(t)$ can be directly counted through detectors at the intersections forming the perimeter border. The authors also claim $q_{i}(t)$ can be available with some applications, such as a substantial percentage of GPS information being readily accessible. However, due to the nature of the prediction horizon in MPC-MHE, such data has to be made available in both real-time and the near future, rendering it impossible to collect, especially under disruptive conditions. We will further tackle such observability issues with the proposed antifragile perimeter control algorithm in the following Section 4.3.
4.2 The DDPG RL algorithm
@@
In RL algorithms, an agent or multiple agents interact with a preset environment and improve its decision-making capacilities, defined as action $a_{k}$ in an action space $\mathcal{A}$, based on the observable state $s_{k}$ in the state space $\mathcal{S}$ and the reward, defined as $r_{k}=\mathcal{R}(s_{k},a_{k})$, where $\mathcal{R}$ is the reward function. The improvement of decision-making is commonly realized through a deep neural network as a function approximator. The RL algorithm applied in this work is Deep Deterministic Policy Gradient (DDPG), as proposed in Lillicrap et al. (2015). By applying an actor-critic scheme, DDPG can manage a continuous action space instead of a discrete action space. Furthermore, Zhou and Gayah (2021) has demonstrated that an RL algorithm with a continuous action space can achieve superior performance. The DDPG algorithm can be divided into two main components, namely the actor and the critic, which are updated at each step through policy gradient and Q-value, respectively. The scheme of the DDPG algorithm applied in this paper is schematically illustrated in Fig. 3.
Figure 3: DDPG scheme.
@@
The actor-network is represented by $\mu(\cdot)$, and it determines the best action $a_{k}$ for the perimeter controller based on the current state $s_{k}$. The critic network $Q(\cdot)$ is responsible for evaluating whether a specific state-action pair at time step $k$ yields the maximal possible discounted future reward $Q(s_{k},a_{k})$. A common technique used in DDPG is to create a target actor network $\mu^{\prime}(\cdot)$ and a target critic network $Q^{\prime}(\cdot)$, which are a copy of the original actor and critic network but updated posteriorly to stabilize the training process and prevent overfitting, with the target maximal discounted future reward calculated as:
$y_{k}=r_{k}+\gamma Q^{\prime}(s_{k+1},\mu^{\prime}(s_{k+1}|\theta^{\mu^{\prime}})|\theta^{Q^{\prime}})$ (5)
@@
The critic network can then be updated by computing the temporal difference between the predicted reward as well as the target reward and minimizing the loss for a mini-batch $N$ sampled from the replay buffer:
$L=\frac{1}{N}\sum_{k}(y_{k}-Q(s_{k},a_{k}|\theta^{Q}))^{2}$ (6)
@@
$I=$ $\mathbb{E}_{s_{k}\sim\rho^{\beta}}[r(s,\mu(s|\theta^{\mu}))|_{s=s_{k}}]$ (7a)
$\nabla_{\theta^{\mu}}I=$ $\mathbb{E}_{s_{k}\sim\rho^{\beta}}[\nabla_{a}Q(s,a|\theta^{Q})|_{s=s_{k},a=\mu(s_{k})}\nabla_{\theta_{\mu}}\mu(s|\theta^{\mu})|_{s=s_{k}}]$ (7b)
$\nabla_{\theta^{\mu}}I\approx$ $\frac{1}{N}\sum_{k}\nabla_{a}Q(s,a|\theta^{Q})|_{s=s_{k},a=\mu(s_{k})}\nabla_{\theta^{\mu}}\mu(s|\theta^{\mu})|_{s_{k}}$ (7c)
First developed in Horgan et al. (2018), a similar Ape-X architecture as applied in Zhou and Gayah (2021) is also adopted in this work, which allows for multiple generations of simulations during training to be included and centrally learned for the best policy. A decaying Gaussian noise is also introduced to encourage the DDPG agent to explore the state space. Starting from a certain training episode, we add disruptions into the simulation environment, such as surging traffic demand or MFD disruption. Such volatility poses additional challenges and opportunities for the learning and decision-making from the agent, leading to updates of its parameters and the capability to yield proper responses after sensing a change in a high-volatility environment.
@@ 4.3 Antifragility and the antifragile terms in RL
As reviewed in Section. 2.2, proper information can be integrated into the RL algorithms to induce robustness or resilience. Following the same idea, we incorporate antifragile derivatives and redundancy terms based on the methodologies reviewed in Section 2.3. First, the benchmark work in Zhou and Gayah (2021) establishes the state as $s_{k}=[n_{ij,k},q_{ij,k}]$, with $q_{ij,k}$ regarded as an estimate of average daily traffic demand. However, real-time traffic demand $q_{ij,k}$ can hardly be acquired, particularly when faced with disruptions. Therefore, we propose a state representation incorporating both the first and second order derivatives of vehicle accumulation, which can be computed as the first and second order differences from the vehicle accumulation $\Delta n_{ij,k},\Delta^{2}n_{ij,k}]$ between two consecutive time steps in the discrete form. Similar to Sirmatel and Geroliminis (2020) discussed in Section 4.1, we also tackle the data observability issue in this work, by integrating trip completion $M_{ij,k}$ and replacing OD-pair accumulation $n_{ij,k}$ with simply regional accumulation $n_{i,k}$ as well we its first and second differences. It should be noted that $M_{ij,k}$ in Sirmatel and Geroliminis (2020) refers specifically to the interregional transfer flow $M_{ij,k},(i\neq j)$. However, when considering a rather homogenous average trip length, the regional trip completion rate $M_{i}(t)$ is linearly related to the network average flow and can be approximated with ease, as illustrated in (Geroliminis and Daganzo, 2008; Zhou and Gayah, 2021). Then the intraregional completion rate can be further computed as $M_{ii,k}=M_{i,k}-M_{ij,k},(i\neq j)$. For fair comparison, we offer two sets of states for testing under either idealized full observability or real-world limited observability, which are studied in Section 6.1 and 6.2, respectively.
@@
The action $a_{k}\in\mathcal{A}$ is defined the same as the control variables $u_{ij}$ in Section 3. For the reward term $r_{k}$, while Zhou and Gayah (2021) uses merely the completion rate, in our proposed algorithm, the reward is defined with additional two terms, i.e., the damping term $r_{dam,k}$ and the redundancy term $r_{red,k}$, as Eq. 4 shows. When the first and second differences $\Delta n_{ij,k},\Delta^{2}n_{ij,k}]$ or $\Delta n_{i,k},\Delta^{2}n_{i,k}]$ are incorporated into the state space $\mathcal{S}$ of the algorithm, significant oscillations in the perimeter control variables can often be observed, especially under scenarios with disruptions. As perimeter control is composed of coordinated traffic lights at the border of different regions, oscillating actions will result in fast varying green splits between consecutive cycles. Even though such oscillations may have a minor impact on traffic performance, the operation of signal lights in the real world should be as stable as possible. Therefore, a damping term $r_{dam,k}$ is introduced into the reward function to penalize potential oscillatory actions. While the absolute difference between two consecutive actions is always below $1$, the constant $\xi_{1}$ bounds the maximal penalty, whereas $\xi_{2}$ determines how the penalty attenuates when the change of control variables becomes small, indicating a large $\xi_{2}$ only penalizes the agent when the oscillation is substantial.
$r_{dam,k}=-\xi_{1}\sum_{i,j\in\{1,2\}}|u_{ij,k}-u_{ij,k-1}|^{\xi_{2}},\quad i\neq j\text{ and }k\neq 1$ (10)
@@
The last term $r_{red,k}$ in the objective function $J_{anti}$ in Eq. 4 acts as an additional term to build up redundancy in the system, resembling the buffer times in train scheduling (Corman et al., 2018). However, instead of a fixed buffer time commonly seen among most European railway networks (Jovanović et al., 2017), we create a dynamic redundancy term emphasizing the derivatives in a similar way as the differences of the regional accumulation, considering real-world limited observability. Here, we summarize $r_{red,k}$ as the sum of two terms $r_{red,k}=H_{k}+\Delta H_{k}$, with $H_{k}$ being an overall term representing the first derivative and $\Delta H_{k}$ repr* esenting the second derivative, which can be further expanded as:
$H_{k}=$ $\sum_{i=1,2}H_{i,k}=\sum_{i=1,2}\omega_{h}\cdot h_{i,k}\cdot\alpha_{i,k}\cdot f(n_{i,k},n_{i,\rm crit},n_{i,\rm cap})$ (11a)
$\Delta H_{k}=$ $\sum_{i=1,2}\Delta H_{i,k}=\sum_{i=1,2}\omega_{\Delta h}\cdot\Delta h_{i,k}\cdot f(n_{i,k},n_{i,\rm crit},n_{i,\rm cap})$ (11b)
@@
In Eq. 11, $\omega_{h}$ and $\omega_{\Delta h}$ are introduced as the weight constants for the derivatives to regulate their impact on the reward $\mathcal{R}$. $h_{i,k}$ and $\Delta h_{i,k}$ are the first and second difference of the traffic states on the MFD, $h_{i,k}$ is defined as the difference of trip completion over vehicle accumulation at the end of a time step versus at the beginning of the same time step, and the second derivative $\Delta h_{i,k}$ is calculated as the difference between the first differences of two consecutive time steps:
$h_{i,k}=\frac{M_{i,k}\ -\ M_{i,k-1}}{n_{i,k}\ -\ n_{i,k-1}}$ (12a)
$\Delta h_{k}=h_{k}-h_{k-1}$ (12b)
@@
The binary variable $\alpha_{i,k}$ in $H_{k}$ was designed to reward the agent when moving towards the desired direction on the MFD. For instance, when the traffic state lies in the congested zone of the MFD, the gradient of any step will be negative. However, a penalty should be applied when vehicle accumulation gets larger, while rewarding the agent when accumulation decreases.
$\alpha_{i,k}=\begin{cases}1,&\text{if }n_{i,k}\geq n_{i,k-1},\\ -1,&\text{otherwise.}\end{cases}$ (13)
@@
Here, $f(n_{i,k},n_{i,\rm crit},n_{i,\rm cap})$ computes a reduction factor to constrain the impact of the $r_{red,k}$ term when the accumulation is around the critical accumulation $n_{i,\rm crit}$, where the $r_{red,k}$ term should have the greatest impact. A modified trigonometric function is applied to realize this purpose. Other functions, such as normal distribution, should also be valid for achieving the same goal.
$f(n_{i,k},n_{i,\rm crit},n_{i,\rm cap})=\begin{cases}{\Big(1+\cos\Big(-\pi\cdot\cfrac{n_{i,\rm crit}-n_{i,k}}{n_{i,\rm crit}}\Big)\Big)}/2,&\text{if }n_{i,k}\geq n_{i,\rm crit},\\ {\Big(1+\cos\Big(-\pi\cdot\cfrac{n_{i,k}-n_{i,\rm crit}}{n_{i,\rm cap}-n_{i,\rm crit}}\Big)\Big)/2},&\text{otherwise.}\end{cases}$ (14)
@@
We illustrate the impact of $H_{k}$ and $\Delta H_{k}$ on building redundancy on the MFD in Fig. 4 and 5. The first derivative $H_{k}$ in Fig. 4(a) rewards the agent when moving toward the critical accumulation $n_{i,\rm crit}$ to maximize trip completion. However, when $n_{i,k}$ approaches $n_{i,\rm crit}$, this term drops significantly and becomes a penalty when $n_{i,k}$ exceeds $n_{i,\rm crit}$. In Fig. 4, we showcase the influence of this term on the MFD. With increasing weight coefficient $\omega_{h}$, $n_{i,\rm crit}$ of the modified MFD becomes lower compared to the original, and the reward decreases faster after the accumulation exceeds $n_{i,\rm crit}$. In this way, redundant overcompensation has been established to prevent accumulation from exceeding the critical accumulation when disruption happens unexpectedly.